OpenStudy (anonymous):

so I have three equations with x, y and z and I need to solve for x, y and z. The first equation is -x+y+4z The second is x+3y+2z=4 The third is 2x+2y-3z=-7 How would I solve this problem?

OpenStudy (anonymous):

it so simple boy use the 3by3 matrix to solve it

OpenStudy (anonymous):

I came up with z = 13, but I can't figure out how to solve for x or y

OpenStudy (anonymous):

do u know how to use echelon form of matrix its really helping when u have to find x y z

OpenStudy (anonymous):

put z=13 in eq2 you will get 2x+2y-39=-7 x=32-2y/2 , z=13 now put values of x, z in any eq it will give value of y then put y in x=32-2y/2 you'll get x y z